# Analytic Model of Radial Sensitivity in Cylindrical PET Systems Based on First Principles

### v2

@@ I-D Computing Solid Angle
Figure 2: Diagram showing the projection of Cap 2 (left cap) onto Cap 1 (right cap) through the point $P$, to form Cap $2^{\prime}$
The projected Cap $2^{\prime}$ will have radius $R_{2}$: $R_{2}=h/(2L-h),$ The center of Cap $2^{\prime}$ will be a distance $m$ from the center of Cap 1, where: $m=r\cdot 2L/(2L-h).$ The total solid angle $\Omega$ is the union of the solid angles subtended by Cap 1 ($\Omega_{1}$) and Cap $2^{\prime}$ ($\Omega_{2}^{\prime}$): $\Omega=\Omega_{1}\cup\Omega_{2}^{\prime}.$
Furthermore, due to the circular symmetry of the system within the XY-plane, it is sufficient to consider radial offsets along the x-axis only. This allows us to disregard the y-coordinate and simplify the source position to the point $P(r,0,L-h).$
@@
$\Omega=\iint\limits_{H}\frac{hdA}{\rho^{3}}=\iint\limits_{H}\frac{hdydx}{[(x-r)^{2}+y^{2}+h^{2}]^{3/2}}$ (7)
@@
$\int\frac{\mathrm{d}x}{(x^{2}+k^{2})^{\frac{3}{2}}}=\frac{x}{k^{2}\cdot\sqrt{x^{2}+k^{2}}}+C$ (8)
@@
$\Omega$ $=\iint\limits_{H}\frac{-h\mathrm{d}y\mathrm{d}x}{[(x-r)^{2}+y^{2}+h^{2}]^{3/2}}$ (9)
@@
$S_{\mathrm{g}}=1-\Omega/2\pi.$ (10)
II Theoretical Analysis and Derivation
II-A Case Analysis
@@
Figure 3: Diagram displaying the case when Cap 1 and Cap $2^{\prime}$ overlap. The blue arc of Cap 1 is denoted as $C_{1}$ and the red arc of Cap $2^{\prime}$ is denoted as $C_{2}$
II-A1 Case 1
In this case, Cap $2^{\prime}$subseteq Cap 1, $\partial H=x^{2}+y^{2}=R^{2}$. We can facilitate the integral by transforming $\partial H$ into parametric form.
@@ II-A1 Case 1
$\Omega$ $=\int_{0}^{2\pi}\frac{h}{[(\cos\theta-r)^{2}+\sin^{2}\theta+h^{2}]^{1/2}}$ (11)
@@
$S_{\mathrm{g}}$ $=1-\frac{1}{2\pi}\int_{0}^{2\pi}\frac{h}{[(\cos\theta-r)^{2}+\sin^{2}\theta+h^{2}]^{1/2}}$ (12)
@@ II-A2 Case 2
The X-coordinate, denoted by $X_{c}$, of the common chord of two intersecting circles is giving by:
$X_{c}=(1-R_{2}^{2}+m^{2})/2m$ (13)
@@
$\Omega$ $=\int_{C_{1}}\frac{-hy\mathrm{d}x}{[(x-r)^{2}+h^{2}]\cdot[(x-r)^{2}+y^{2}+h^{2}]^{1/2}}$ (14)
@@
$S_{\mathrm{g}}$ $=1-\frac{\Omega}{2\pi}$ (15)
@@
II-B Theoretical Result
We consolidate the two cases into a single result, where the general geometric sensitivity of the system is modeled by the following distribution:
$S_{\mathrm{g}}=\begin{cases}\text{equation (12),}&\text{$m+R_{2}\leq R$}\\ \text{equation (15),}&\text{$m+R_{2}>R$}\\ \end{cases}$ (16)
Using MATLAB[22], we plot the sensitivity $S_{\mathrm{g}}$ based on equation (16) with 100 evenly spaced points in the range [0, R]. The dimensions of the PET system were set to $R=405$ mm and $L=153$ mm. The resulting plot is shown in Figure 4.
Figure 4: Plot of the geometric sensitivity as a function of radial offset, calculated based on equation (16). The plot considers 100 evenly spaced points within $[0,R]$ for a PET system with dimensions $R=405$ mm and $L=153$ mm
III Experiments and Simulations
(a) Simulation with FOV=250mm
(b) Simulation with FOV=300mm
Figure 5: Distributions for a) sensitivity w.r.t radial offset for a FOV radius of $250$ mm across 4 axial slices, and b) sensitivity w.r.t radial offset for a FOV radius of $300$ mm across 4 axial slices
III-A GATE Simulations
Simulations were performed in GATE[23] using a PET scanner with a radius $R=403.8$ mm and an axial length $3062L=$ 306 mm. The scanner used lutetium yttrium osmium (LYSO) scintillator crystals, arranged in a $6\times 6$ grid within each detector block. Each crystal had dimensions of $20.0\,\text{mm}\times 3.9\,\text{mm}\times 3.9\,\text{mm}$.
A cylindrical ${}^{18}\text{F}$ radioactive source with a radius of $0.01$ mm and a length of $700$ mm was used. The source had an activity of $7.4$ MBq ($7.4\times 10^{6}$ Bq).
The following thresholding methods were applied:
  • Energy Window Filtering: Only events with energy levels within the window of $[420,650]\,\text{keV}$ were considered valid.
  • Distance Thresholding/FOV: Events detected outside the FOV radius were discarded. Simulations were conducted for 250 mm FOV radius and 300 mm FOV radius.
Sensitivity values were obtained at radial offsets in increments of $25$ mm, covering a range from 00 mm to $400$ mm, and at axial slices in increments of $25$ mm, covering a range from 00 mm to $75$ mm[24]. These values were calculated in accordance with the NEMA standard [8]. The corresponding plots are shown in Figure 5.
III-B Experiments
Experiments were performed using the same PET scanner dimensions as in the simulations, with a ring radius of $403.8$ mm and an axial length of $306$ mm.
A cylindrical ${}^{18}\text{F}$ radioactive source with a radius of $0.01$ mm and a length of $700$ mm was used, with an activity of $7.4$ MBq. Data collection and sensitivity calculations followed the NEMA standard[8], ensuring comparability with established benchmarks.
Figure 6: Experimental sensitivity across axial slices for 16 different radial offsets
Figure 7: Experimental sensitivity w.r.t radial offset. The sensitivity measurements were taken at $25$ mm intervals from 00 mm to $375$ mm, resulting in 16 distinct points plotted in the figure
The following thresholding methods were applied:
  • Energy Window Filtering: Only events with energy levels within the window of $[420,650]\,\text{keV}$ were considered valid.
  • Distance Thresholding/FOV: A 250 mm FOV radius was enforced, discarding events detected beyond this range.
Sensitivity values were measured after subtraction of random coincidences. Figure 6 presents the sensitivity distribution across 150 axial slices at 16 radial offsets ranging from 00 mm to $375$ mm[24]. Figure 7 illustrates the sensitivity distribution with respect to radial offset at the center slice of the PET scanner.
IV Discussion
(a) Experimental Result
(b) Theoretical Result
(c) Simulation With FOV=250mm
(d) Simulation With FOV=300mm
Figure 8: Side-by-side compilation of previous diagrams of experimental, theoretical, and simulated sensitivity for direct comparison
The theoretical model derived in this study predicts that system sensitivity in a cylindrical PET scanner first slightly decreases then increases with radial offset and reaches a maximum near the periphery of the FOV. To evaluate the applicability of this model, we compared its predictions against results obtained from simulations and physical measurements. The results are put side-by-side in Figure 8.
The simulation results largely mirror the theoretical curve, confirming the expected increase in sensitivity with radial offset and locating the peak near the outer radial edge of the scanner’s FOV. Empirical results, while consistent with this general trend, show noticeable differences. In particular, both the simulation and experimental sensitivity distribution exhibit a sharp drop off past a certain radial offset. These differences are to be expected, given that the experimental data reflect a range of real-world factors that are deliberately excluded from the idealized model.
Despite these deviations, the agreement between the theoretical and empirical trends at a qualitative level is both meaningful and reassuring. The model accurately captures the overall behavior of sensitivity across the radial plane and establishes a reliable reference for identifying and interpreting departures in practical systems. The regions of agreement—especially the general increase in sensitivity away from the center—affirm the validity of the model’s prediction, while the observed differences highlight the influence of non-ideal system characteristics.
Future work will focus on investigating system-level and implementation-specific factors that influence sensitivity, to better understand their contribution to the discrepancies between theoretical predictions and empirical observations. By systematically introducing variables into extended simulations or layered analytic models, we aim to quantify their individual contributions and refine the theoretical description accordingly.
V Conclusion
The analytic model derived in this study provides a first-principles description of the radial-plane sensitivity distribution for cylindrical PET systems under idealized conditions. By isolating the system geometry and basic detection physics, the model establishes a theoretical baseline that had not previously been articulated in closed form. The formulation presented here contributes a conceptual tool for understanding how system sensitivity varies with radial position in an ideal setting.
Beyond its conceptual value, the model can also serve as a practical baseline for system evaluation, scanner calibration, and design analysis. By excluding confounding factors such as detector efficiency, crystal geometry, attenuation, or time-of-flight corrections, the theoretical curve can be used to isolate and study the influence of these variables in real systems. Thus, this work is not intended to refute prior assumptions or findings, but rather to supplement them by providing a precise theoretical framework.
Acknowledgment

### v3

@@ Appendix A Appendix: Computing Solid Angle
Figure 6: Diagram showing the projection of Cap 2 (left cap) onto Cap 1 (right cap) through the point $P$, to form Cap $2^{\prime}$.
The projected Cap $2^{\prime}$ will have radius $R_{2}$: $R_{2}=R(L-h)/(L+h),$ The center of Cap $2^{\prime}$ will be a distance $m$ from the center of Cap 1, where: $m=r\cdot 2L/(2L-h).$ The total solid angle $\Omega$ is the union of the solid angles subtended by Cap 1 ($\Omega_{1}$) and Cap $2^{\prime}$ ($\Omega_{2}^{\prime}$): $\Omega=\Omega_{1}\cup\Omega_{2}^{\prime}.$
Furthermore, due to the circular symmetry of the system within the XY-plane, it is sufficient to consider trans-axial offsets along the x-axis only. This allows us to disregard the y-coordinate and simplify the source position to the point P(r,0,Lh).P(r,0,L-h).
@@
$\Omega=\iint\limits_{H}\frac{hdA}{\rho^{3}}=\iint\limits_{H}\frac{hdydx}{[(x-r)^{2}+y^{2}+h^{2}]^{3/2}}$ (A.3)
@@
$\int\frac{\mathrm{d}x}{(x^{2}+k^{2})^{\frac{3}{2}}}=\frac{x}{k^{2}\cdot\sqrt{x^{2}+k^{2}}}+C$ (A.4)
@@
$\Omega$ $=\iint\limits_{H}\frac{-h\mathrm{d}y\mathrm{d}x}{[(x-r)^{2}+y^{2}+h^{2}]^{3/2}}$ (A.5)
@@
$S_{\mathrm{g}}=1-\Omega/2\pi.$ (A.6)
Appendix B Appendix: Case Analysis
@@
Figure 7: Diagram displaying the case when Cap 1 and Cap $2^{\prime}$ overlap. The blue arc of Cap 1 is denoted as $C_{1}$ and the red arc of Cap $2^{\prime}$ is denoted as $C_{2}$.
B.1 Case 1
In this case, Cap $2^{\prime}$ $\subseteq$ Cap 1, $\partial H=x^{2}+y^{2}=R^{2}$. We can facilitate the integral by transforming $\partial H$ into parametric form.
@@ B.1 Case 1
$\Omega$ $=\int_{0}^{2\pi}\frac{h}{[(\cos\theta-r)^{2}+\sin^{2}\theta+h^{2}]^{1/2}}$ (B.1)
@@
$S_{\mathrm{g}}$ $=1-\frac{1}{2\pi}\int_{0}^{2\pi}\frac{h}{[(\cos\theta-r)^{2}+\sin^{2}\theta+h^{2}]^{1/2}}$ (B.2)
@@ B.2 Case 2
The X-coordinate, denoted by $X_{c}$, of the common chord of two intersecting circles is given by:
$X_{c}=(1-R_{2}^{2}+m^{2})/2m$ (B.3)
@@
$\Omega$ $=\int_{C_{1}}\frac{-hy\mathrm{d}x}{[(x-r)^{2}+h^{2}]\cdot[(x-r)^{2}+y^{2}+h^{2}]^{1/2}}$ (B.4)
@@
$S_{\mathrm{g}}$ $=1-\frac{\Omega}{2\pi}$ (B.5)